# Why Knowledge Distillation Works in Generative Models: A Minimal Working Explanation

### v1

@@ 4.1 From Gaussian Mixtures to Autoregressive Language Models
While the previous section focused on synthetic data from Gaussian mixtures, the core mechanism we observed—namely, a trade-off between precision and recall modulated by teacher entropy—extends naturally to autoregressive language models. Distillation has long been a standard technique in training such models (kim2016sequence, ), and is now integral to large-scale language model (LLM) pipelines, from instruction tuning to efficient deployment in resource-constrained environments (ko2024distillm, ; abouelenin2025phi, ; grattafiori2024llama, ).
@@
mirroring the mixture structure seen in Gaussian models. Moreover, it is common practice to apply entropy-reducing techniques—such as temperature scaling, top-$k$, or top-$p$ sampling (welleck-etal-2020-consistency, )—to produce sharper generations. This is analogous to our $\beta$-modulated teacher distributions.
Prior work (yang2018breaking, ) further shows that the expressivity of each token distribution is upper-bounded by the dimensionality of the model’s hidden states. Since this dimensionality scales with model size, smaller models inherently represent fewer modes, and thus face similar capacity bottlenecks to those seen in our mixture of Gaussian setting.
In the following section, we test whether the precision–recall trade-off observed in simulation also emerges in LLMs. We perform multistage distillation using the SmolLM2 (allal2025smollm2, ) family of models, and examine how increasing teacher selectivity—controlled via a $\beta$-like parameter—affects the student’s precision and recall.
@@ 4.2 Experimental setup
To test whether the precision–recall dynamics observed in our synthetic setting also manifest in real-world models, we conduct a series of distillation experiments using the SmolLM2 (allal2025smollm2, ) family. Fig. 3 provides an overview of our experimental pipeline, which mirrors the generative structure outlined in Section 3.
@@
To construct the teacher model $p^{\prime}$ (Equation 1), we randomly initialize a 360M parameter SmolLM2 model and pretrain it from scratch on $D$ using the standard next-token prediction loss. Training is conducted using 4 GPUs with a total batch size of 256, optimized using AdamW loshchilov2017decoupled (initial learning rate 5e–4) with a warmup–square-decay scheduler zhai2022scaling . We train for 5 epochs and follow hyperparameter settings from the original SmolLM2 release, with minor adjustments for training stability as detailed in the Supplementary. We select the best-performing teacher checkpoint based on the lowest perplexity evaluated on a held-out validation set $D_{\text{val}}$ of 100,000 sentences sampled independently from $p^{*}$.
Next, we distill this teacher model into a 135M student model $p^{\prime\prime}$ (in Equation 2) using temperature-scaled generations from $p^{\prime}$. Specifically, we sample from the pretrained teacher $p^{\prime}$ using multiple temperatures ($\tau\in\{0.8,0.875,0.95,1.0\}$), which correspond to varying levels of entropy in the output distribution—lower temperatures produce more selective (i.e., lower-entropy) teachers. For each temperature, we generate a new dataset $D^{\prime}_{\tau}$, consisting of 10M sequences, and train the student model $p^{\prime\prime}$ from scratch using the same training setup and hyperparameters as for the teacher.
Finally, we measure the student model’s precision and recall with respect to the original 1.7B model $p^{*}$ using the definitions in Equations 3 and 4. Precision is computed using 100,000 validation sequences ($D^{\prime\prime}_{\text{val},\tau}$) sampled with top-k using temperature $\tau=1.0$ from the student model $p^{\prime\prime}(D^{\prime}_{\tau})$ , while recall is measured using the same number of samples from the ground-truth $p^{*}$ (i.e., from $D_{\text{val}}$). This evaluation allows us to observe how the precision–recall trade-off evolves as the teacher becomes more selective via temperature scaling.
@@ 4.3 Experimental result
Models $p^{\prime\prime}(D^{\prime}_{\tau=0.8})$ $p^{\prime\prime}(D^{\prime}_{\tau=0.875})$ $p^{\prime\prime}(D^{\prime}_{\tau=0.95})$ $p^{\prime\prime}(D^{\prime}_{\tau=1.0})$ $p^{\prime\prime}(D)$
Precision
Recall
-0.7893
-4.3469
-2.2418
-4.2420
-3.7284
-4.0858
-4.5213
-4.0453
-5.5957
-4.0122
Table 1: Comparison of precision and recall on trained models. Higher (less negative) values indicate better precision/recall.
Table 1 reports the precision and recall of student models $p^{\prime\prime}$ (SmolLLM2 135M) distilled from the teacher $p^{\prime}$ (SmolLLM2 360M) under varying sampling temperatures $\tau$, with evaluation performed against the ground-truth model $p^{*}$ (pretrained SmolLLM2 1.7B). Each column $p^{\prime\prime}(D^{\prime}_{\tau})$ corresponds to a specific distillation setup, where $p^{\prime\prime}$ is trained on samples drawn from $p^{\prime}$ using temperature $\tau$. The final column, $p^{\prime\prime}(D)$, represents a baseline where the student is trained directly on data sampled from $p^{*}$, without any intermediate teacher.
We observe that as the sampling temperature $\tau$ decreases, the student model $p^{\prime\prime}$ exhibits higher precision but lower recall—favoring high-likelihood regions while sacrificing coverage. This behavior suggests that lower temperatures lead the teacher to produce more peaked outputs, concentrating mass on dense modes, which the student mimics—trading off coverage for sharpness. The result closely mirrors our earlier findings in the Gaussian mixture setting, where distillation from a low-entropy teacher led the student to replicate only the most prominent modes of the original distribution.
These results provide empirical evidence that the geometric behavior of knowledge distillation—trading off recall for increased selectivity—persists even in large-scale language models. Rather than merely compressing the teacher’s output space, distillation reshapes the student’s distribution to emphasize its high-density core. This effect is especially useful in scenarios where sample quality is more critical than diversity, such as instruction tuning or downstream generation tasks. Our findings thus reinforce the interpretation of distillation as a mechanism for inducing controlled distributional concentration, consistent with the minimal setting analyzed earlier.
(a) $\tau=0.8$
@@ 4.4 Embedding Space Visualization
To further examine how distillation affects the generative behavior of student models, we visualize the distribution of generated samples from $p^{*}$ and $p^{\prime\prime}(D^{\prime}_{\tau})$ in a shared embedding space using Nomic Embed v2 (nussbaum2025trainingsparsemixtureexperts, ). While the true distribution $p^{*}$ in language space cannot be directly plotted, sentence embeddings provide a meaningful proxy for analyzing geometric properties of generated text. Fig. 4 presents 2D projections (via UMAP (mcinnes2018umap, )) of sampled sentences from two sources: the ground-truth model $p^{*}$ (gray) and student models trained via distillation from the teacher using different sampling temperatures (red). Each subplot corresponds to a different value of $\tau$ used during teacher sampling.
@@
These results confirm that the distributional narrowing induced by distillation—previously observed in the Gaussian mixture setup and through quantitative precision–recall metrics—also manifests geometrically in large-scale language models. While embedding projections provide only a coarse approximation of semantic structure, they visually reinforce our interpretation of distillation as inducing controlled concentration in the student’s output distribution.
@@ 5 Concluding Remarks
A principled explanation for knowledge distillation in generative modeling.   Despite its widespread adoption in modern LLM pipelines, knowledge distillation remains largely treated as a heuristic—a black-box technique assumed to work without a concrete understanding of how or why. In this paper, we presented what we believe to be the first principled analysis of how distillation affects generative behavior, both in theory and in practice. Starting from a controlled mixture-of-Gaussians setup, we demonstrated that distillation naturally induces a precision–recall trade-off governed by the entropy of the teacher distribution. Crucially, this theoretical behavior was not only consistent across simulations but also emerged clearly in experiments with autoregressive language models.
@@
A minimal yet important working explanation.   By systematically varying the selectivity of the teacher and quantifying the resulting student behaviors, we provided a minimal working explanation of what knowledge distillation actually does in the generative context: it reshapes the student’s distribution to concentrate on high-probability regions prioritized by the teacher. This effect is especially pronounced when the teacher is low in entropy (e.g., via temperature scaling), leading to sharper but less diverse outputs. When sample quality is preferred over full coverage—a common scenario in instruction tuning and downstream generation—this trade-off becomes not just acceptable, but desirable. We believe our framework fills a critical gap in the literature by offering a geometric, distribution-level interpretation of distillation applicable to both toy and large-scale generative models.
@@
Limitations and future directions.   While our analysis offers a foundational step toward demystifying knowledge distillation, several limitations remain. Most notably, our LLM experiments focus on the pretraining phase—distilling from synthetic teacher data generated via next-token prediction. However, in practice, distillation is also widely applied during post-training stages, such as instruction tuning, alignment, or preference modeling. Future work should examine whether the same precision–recall trade-off persists in these settings, and how it interacts with additional fine-tuning signals. We anticipate that the underlying dynamics observed in this study will remain valid, but merit further validation in applied settings.
@@
Societal impact.   Our findings suggest that distillation can improve generation efficiency and reduce deployment cost, contributing positively to accessibility and energy consumption. However, compressing generative capabilities into smaller models may also lower barriers to misuse (e.g., spam or disinformation), warranting careful access control and monitoring during deployment.
@@
In summary, we view this work as a step toward grounding the use of distillation in generative modeling in theoretical and empirical understanding. By identifying and validating a minimal working explanation, we hope to shift distillation from a rule of thumb to a principled design tool—one that can be better tuned, adapted, and trusted in the development of future generative model design.
6 Acknowledgement
@@ A.1 Pretraining Setup
$p^{\prime}(D)$ is trained for 5 epochs using 4 NVIDIA V100 GPUs, with DeepSpeed rasley2020deepspeed ZeRO Stage 2 optimization and FP16 mixed-precision enabled. The global batch size is 256, achieved using a per-device micro-batch size of 32 and a gradient accumulation step size of 2. We use the AdamW optimizer loshchilov2017decoupled with a learning rate of $5\times 10^{-4}$, $(\beta_{1},\beta_{2})=(0.9,0.95)$, and no weight decay. We apply a WSD learning rate schedule zhai2022scaling : the first 1% of training steps are reserved for linear warmup, followed by a stable learning rate phase, and a final linear decay over the last 20% of steps. $p^{\prime\prime}(D^{\prime}_{\tau})$ is trained with the same settings, except it is only trained for one epoch. All models are initialized from the HuggingFace checkpoints HuggingFaceTB/SmolLM2-{size}M, and their weights are fully reinitialized prior to training.
@@ A.2 UMAP-Based Visualization Setup
To analyze the distributional characteristics of the learned models, we project sentence embeddings into a 2D space using UMAP mcinnes2018umap . For each model distribution (e.g., $p^{*}$, $p^{\prime\prime}(D^{\prime}_{\tau})$), we randomly sample 4,000 responses from $p^{*}$ and 1,000 responses from each $p^{\prime\prime}(D^{\prime}_{\tau})$, and compute their embeddings using the all-mpnet-base-v2 model provided by Nomic Embed v2 (nussbaum2025trainingsparsemixtureexperts, ). The UMAP projection is configured with n_neighbors=5, min_dist=0.001, and metric=‘cosine’.

### v3

@@ 4.1 From Gaussian mixtures to autoregressive language models
While the previous section focused on synthetic data from Gaussian mixtures, the core mechanism we observed—namely, a trade-off between precision and recall modulated by teacher entropy—extends naturally to autoregressive language models. Distillation has long been a standard technique in training such models (Kim and Rush, 2016), and is now integral to large-scale language model (LLM) pipelines, from instruction tuning to efficient deployment in resource-constrained environments (Ko et al., 2024; Abouelenin et al., 2025; Grattafiori et al., 2024).
@@
mirroring the mixture structure seen in Gaussian models. Moreover, it is common practice to apply entropy-reducing techniques, such as temperature scaling, top-$k$, or top-$p$ sampling (Welleck et al., 2020), to produce sharper generations. This is analogous to our $\beta$-modulated teacher distributions.
Prior work (Yang et al., 2018) further shows that the expressivity of each token distribution is upper-bounded by the dimensionality of the model’s hidden states. Since this dimensionality scales with model size, smaller models inherently represent fewer modes, and thus face similar capacity bottlenecks to those seen in our mixture of Gaussian setup.
In the following section, we test whether the precision-recall trade-off observed in simulation also emerges in LLMs. We perform multistage distillation using the SmolLM2 (Allal et al., 2025) family of models, and examine how increasing teacher selectivity (controlled via a $\beta$-like parameter) affects the student’s precision and recall.
@@ 4.2 Experimental setup
To test whether the precision-recall dynamics observed in our synthetic setup also manifest in real-world models, we conduct a series of distillation experiments using the SmolLM2 (Allal et al., 2025) family. Fig. 3 provides an overview of our experimental pipeline, which mirrors the generative structure outlined in Section 3. We conduct each experiment with five different random seeds.
@@
To construct the teacher model $p^{\prime}$ (Equation 1), we randomly initialize a 360M parameter SmolLM2 model and pretrain it from scratch on $D$ using the standard next-token prediction loss. Training is conducted using 4 GPUs with a total batch size of 256, optimized using AdamW Loshchilov and Hutter (2017) (initial learning rate 5e-4) with a warmup-square-decay scheduler Zhai et al. (2022). We train for 5 epochs and follow hyperparameter settings from the original SmolLM2 release, with minor adjustments for training stability as detailed in the Supplementary. We select the best-performing teacher checkpoint based on the lowest perplexity evaluated on a held-out validation set $D_{\text{val}}$ of 100,000 sentences sampled independently from $p^{*}$.
Next, we distill this teacher model into a 135M student model $p^{\prime\prime}$ (in Equation 2) using temperature-scaled generations from $p^{\prime}$. Specifically, we sample from the pretrained teacher $p^{\prime}$ using multiple temperatures ($\tau\in\{0.8,0.875,0.95,1.0\}$), which correspond to varying levels of entropy in the output distribution—lower temperatures produce more selective (i.e., lower-entropy) teachers. For each temperature $\tau$, we generate a new dataset $D^{\prime}_{\tau}$ consisting of 10M sequences. We then train the student model $p^{\prime\prime}$ from scratch on each $D^{\prime}_{\tau}$ using the same training setup and hyperparameters as for the teacher.
Finally, we measure the student model’s precision and recall with respect to the original 1.7B model $p^{*}$ using the definitions in Equations 3 and 4. Precision is computed using 100,000 validation sequences ($D^{\prime\prime}_{\text{val},\tau}$) sampled with top-k using temperature $\tau=1.0$ from the student model $p^{\prime\prime}(D^{\prime}_{\tau})$ , while recall is measured using the same number of samples from the ground-truth $p^{*}$ (i.e., from $D_{\text{val}}$). This evaluation allows us to observe how the precision-recall trade-off evolves as the teacher becomes more selective via temperature scaling.
@@ 4.3 Experimental result
Figure 4: Score distribution of Precision (left) and Recall (right) based on the $\tau$ parameter and the $P^{\prime\prime}(D)$ model. Each box plot illustrates the interquartile range across five seeds, with the orange line indicating the arithmetic mean. Higher (less negative) values on the y-axis denote better result.
Fig. 4 reports the precision and recall of student models $p^{\prime\prime}$ (SmolLM2 135M) distilled from the teacher $p^{\prime}$ (SmolLM2 360M) under varying sampling temperatures $\tau$, with evaluation performed against the ground-truth model $p^{*}$ (pretrained SmolLM2 1.7B). Each $\tau$ in x-axis corresponds to a specific distillation setup, where $p^{\prime\prime}$ is trained on samples drawn from $p^{\prime}$ using temperature $\tau$ ($p^{\prime\prime}(D^{\prime}_{\tau})$). The final one, $p^{\prime\prime}(D)$, represents a baseline where the student is trained directly on data sampled from $p^{*}$, without any intermediate teacher.
We observe that as the sampling temperature $\tau$ decreases, the student model $p^{\prime\prime}$ exhibits higher precision but lower recall, favoring high-likelihood regions while sacrificing coverage. This behavior suggests that lower temperatures lead the teacher to produce more peaked outputs, concentrating mass on dense modes, which the student mimics, thereby trading off coverage for sharpness. The result closely mirrors our earlier findings in the Gaussian mixture framework, where distillation from a low-entropy teacher led the student to replicate only the most prominent modes of the original distribution.
These results provide empirical evidence that the geometric behavior of knowledge distillation—trading off recall for increased precision—persists even in large-scale language models. Rather than merely compressing the teacher’s output space, distillation reshapes the student’s distribution to emphasize its high-density core. This behavior is especially useful in scenarios where sample quality is more critical than diversity, such as instruction tuning, reasoning tasks, or downstream generation tasks. Our findings thus reinforce the interpretation of distillation as a mechanism for inducing controlled distributional concentration, consistent with the Gaussian mixture framework analyzed earlier.
@@ 4.4 Embedding space visualization
To further examine how distillation affects the generative behavior of student models, we visualize the distribution of generated samples from $p^{*}$ and $p^{\prime\prime}(D^{\prime}_{\tau})$ in a shared embedding space using Nomic Embed v2 (Nussbaum and Duderstadt, 2025). While the true distribution $p^{*}$ in language space cannot be directly plotted, sentence embeddings provide a meaningful proxy for analyzing geometric properties of generated text. Fig. 5 presents 2D projections (via UMAP (McInnes et al., 2018)) of sampled sentences from two sources: the ground-truth model $p^{*}$ (gray) and student models trained via distillation from the teacher using different sampling temperatures (red). Each subplot corresponds to a different value of $\tau$ used during teacher sampling.
@@
These results confirm that the distributional narrowing induced by distillation—previously observed in the Gaussian mixture framework and through quantitative precision-recall metrics—also manifests geometrically in large language models. While embedding projections provide only a coarse approximation of semantic structure, they visually reinforce our interpretation of distillation as inducing controlled concentration in the student’s output distribution.
@@ 5 Concluding Remarks
A principled explanation for knowledge distillation in generative modeling. Despite its widespread adoption in generative models, knowledge distillation remains largely treated as a heuristic: a black-box technique assumed to work without a concrete understanding of how or why. In this paper, we presented what we believe to be the first principled analysis of how distillation affects generative behavior, both in theory and in practice. Starting from a controlled mixture of Gaussians setup, we demonstrated that distillation naturally induces a precision-recall trade-off governed by the entropy of the teacher distribution. Crucially, this theoretical behavior was not only consistent across simulations but also emerged clearly in experiments with autoregressive large language models.
@@
A minimal yet important working explanation. By systematically varying the selectivity of the teacher and quantifying the resulting student behaviors, we provided a minimal working explanation of what knowledge distillation actually does in the generative context: it reshapes the student’s distribution to concentrate on high-probability regions prioritized by the teacher. This effect is especially pronounced when the teacher exhibits low entropy (e.g., via temperature scaling), leading to sharper but less diverse outputs. When sample quality is preferred over full coverage, as is common in instruction tuning, reasoning tasks, and downstream tasks, this trade-off becomes not just acceptable, but desirable. Our framework fills a critical gap in the literature by offering a geometric, distribution-level interpretation of distillation applicable to both toy and large-scale generative models.
@@
Limitations and future directions. While our analysis offers a foundational step toward demystifying knowledge distillation, several limitations remain. Most notably, our LLM experiments focus on the pretraining phase, distilling from synthetic teacher data generated via next-token prediction. However, in practice, distillation is also widely applied during post-training stages, such as instruction tuning, alignment, or preference modeling. Future work should examine whether the same precision-recall trade-off persists in these settings, and how it interacts with additional fine-tuning signals. We anticipate that the underlying dynamics observed in this study will remain valid, yet they merit further validation in applied settings.
@@
Societal impact. Our findings suggest that distillation can improve generation efficiency and reduce deployment cost, contributing to improved accessibility and reduced energy consumption. However, compressing generative capabilities into smaller models may also lower barriers to misuse (e.g., spam or disinformation), warranting careful access control and monitoring during deployment.
@@
In summary, we view this work as a step toward grounding the use of distillation for generative modeling in theoretical and empirical understanding. By identifying and validating a minimal working explanation, we hope to shift distillation from a rule of thumb to a principled design tool—one that can be better tuned, adapted, and relied upon in the development of future generative models.
Acknowledgement
@@ A.1 Pretraining setup
$p^{\prime}(D)$ is trained for 5 epochs using 4 NVIDIA V100 GPUs, with DeepSpeed Rasley et al. (2020) ZeRO Stage 2 optimization and FP16 mixed-precision enabled. The global batch size is 256, achieved using a per-device micro-batch size of 32 and a gradient accumulation step size of 2. We use the AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of $5\times 10^{-4}$, $(\beta_{1},\beta_{2})=(0.9,0.95)$, and no weight decay. We apply a WSD learning rate schedule Zhai et al. (2022): the first 1% of training steps are reserved for linear warmup, followed by a stable learning rate phase, and a final linear decay over the last 20% of steps. $p^{\prime\prime}(D^{\prime}_{\tau})$ is trained with the same settings, except it is only trained for one epoch. All models are initialized from the HuggingFace checkpoints: HuggingFaceTB/SmolLM2-{size}M, and their weights are fully reinitialized prior to training.
@@ A.2 UMAP-Based visualization setup
To analyze the distributional characteristics of the learned models, we project sentence embeddings into a 2D space using UMAP McInnes et al. (2018). For each model distribution (e.g., $p^{*}$, $p^{\prime\prime}(D^{\prime}_{\tau})$), we randomly sample 4,000 responses from $p^{*}$ and 1,000 responses from each $p^{\prime\prime}(D^{\prime}_{\tau})$, and compute their embeddings using the all-mpnet-base-v2 model provided by Nomic Embed v2 (Nussbaum and Duderstadt, 2025). The UMAP projection is configured with n_neighbors=5, min_dist=0.001, and metric=‘cosine’.
Appendix B Additional Discussion
B.1 How the precision-recall trade-off manifests across different downstream tasks
Our framework, which identifies knowledge distillation (KD) as inducing a precision-recall trade-off, can be extended to provide a principled explanation for how distilled models behave across a spectrum of distinct downstream tasks, such as summarization, reasoning, or Chain-of-Thought (CoT) generation.
We can conceptualize the modes in our generative simulation (Fig. 1a) as analogies for the data distributions of these distinct capabilities. In this view, each mode represents a specific skill or task. A large, generalist teacher model must possess broad capabilities, requiring it to cover this entire multi-modal landscape. This corresponds to achieving high recall over the complete set of tasks.
Our work demonstrates that the distillation process, particularly with a selective, lower-entropy teacher (as modulated by $\beta$ in our simulation or a low temperature $\tau$ in language models), guides the student to concentrate its probability mass on a specific subset of these modes (as empirically shown in Fig. 2c and Fig. 5). This results in a student model that achieves high precision: it develops a strong, focused competence on the targeted tasks, often matching or even exceeding the teacher’s performance on those specific skills.
This gain in precision, however, occurs concurrently with a reduction in recall. The student model may lose its generalist abilities or perform poorly on the non-targeted tasks (modes) that were de-emphasized during distillation. This provides a distribution-level explanation for a widely observed empirical phenomenon: distilled student models frequently exhibit a loss of general capability while demonstrating exceptional performance on the specific tasks for which they were optimized.
To offer a concrete example, consider distilling a model for CoT reasoning. Our framework suggests that the resulting student might become highly proficient at generating a particular style or format of reasoning chain (high precision), especially if that style was dominant in the teacher’s selective demonstrations. However, this student might simultaneously lose the ability to generate more diverse, exploratory, or alternative reasoning paths (low recall) that the original generalist teacher was capable of producing.